I need help on this problem: what is the supplement of 125 degrees?
Answer by drj(1380)

I need help on this problem: what is the supplement of 125 degrees?

Step 1. Supplementary angles mean two angles add up to 180 degrees.

Step 2. Given x=125 degrees

Step 3. Then, 180-x=180-125=55 be the supplement of 125 degrees.

Step 4. ANSWER: The supplement of 125 degrees is 55 degrees.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J